# Efficient tau-pair invariant mass reconstruction with simplified matrix element techniques

### v2

@@ Abstract
The quality of the invariant mass reconstruction of the di-$\tau$ system is crucial for searches and analyses of di-$\tau$ resonances. Due to the presence of neutrinos in the final state, the $\tau\tau$ invariant mass cannot be calculated directly at hadron colliders, where the longitudinal momentum sum constraint cannot be applied. A number of approaches have been adopted to mitigate this issue. The most general one uses Matrix Element (ME) integration for likelihood estimation, followed by invariant mass reconstruction as the value maximizing the likelihood. However, this method has a significant computational cost due to the need for integration over the phase space of the decay products. We propose an algorithm that reduces the computational cost by two orders of magnitude, while maintaining the resolution of the invariant mass reconstruction at a level comparable to the ME-based method. Moreover, we introduce additional features to estimate the uncertainty of the reconstructed mass and the kinematics of the initial $\tau$ leptons (e.g., their momenta).
@@ 1 Introduction
The $\tau\tau$ invariant mass can not be calculated from the quantities measured by experiments at hadron colliders, as there are at least two neutrinos in the final state. In the case of neutrinos, the only available information is the total transverse momentum carried out by all neutrinos estimated by the missing transverse energy, denoted $\overline{E}_{T}^{\text{miss}}$. Even though the momenta of the remaining neutral and charged decay products $\tau$ (visible products, called $\tau_{\text{vis}}$) can be precisely measured, there are two main difficulties:
@@
    incomplete description of the final state kinematics;
@@
    resolution $\overline{E}_{T}^{\text{miss}}$, which is one of the least precisely reconstructed observables in a hadronic environment.
Two main approaches to mitigate these issues, adopted by the ATLAS and CMS Collaborations, are based on the maximum likelihood, $\mathcal{L}(m_{\text{test}}|\text{data})$, parameter estimation (MLE), where $m_{\text{test}}$ is the tested $\tau\tau$ invariant mass hypothesis and ”data” is all information available from the detector. However, both collaborations define the likelihood in a different way.
In the case of the ATLAS Collaboration, the likelihood is based on probability density functions (PDFs) of a distance between neutrinos and visible tau decay products [9]. The neutrino direction hypotheses are raised, and the one yielding the maximum likelihood value is chosen. Reconstructed neutrino direction, together with measured visible momenta, allows for a $\tau\tau$ invariant mass calculation. The PDFs are obtained from a large sample of Monte Carlo simulations of $Z\rightarrow\tau\tau$ decays.
The CMS Collaboration adopted the matrix element-based technique in two versions: with the full matrix element for $pp\rightarrow H\rightarrow\tau\tau\rightarrow\dots$ and all relevant normalization factors included (denoted Secondary Vertex fit (SVfit)) and a simplified version (classic Secondary Vertex Fit (cSVfit)) [4]. In the simplified version, only terms corresponding to the $H\rightarrow\tau\tau\rightarrow\dots$ decay are retained. Both methods show similar performance in terms of mass resolution, but the simplified method requires significantly less computational time – around $0.25$ $s$ on a modern computer. However, it is still too much for complex analyses, where a systematic uncertainty estimate requires multiple calculations of the invariant mass for each event, and event samples to process have sizes of the order of $10^{6}$ events.
In this article, we introduce a fast algorithm (fastMTT) inspired by the cSVfit. We merge the idea of cSVfit with the well-known collinear approximation (CA) [10] approach. After applying the CA approximation to the likelihood calculation, we have achieved approximately 100 times speedup in the code execution time, while maintaining comparable mass resolution. The algorithm has been briefly described in the proceedings of the Epiphany 2025 conference [15]. In this paper, we present its full details.
@@ 2 Maximum likelihood estimation
$\mathcal{L}(\text{data}|m_{\text{test}})=$ $\,\mathcal{N}\int d\Phi_{n}\,\left|\mathcal{M}(\text{true}|m_{\text{test}})\right|^{2}W(\text{data}|\text{true}),$ (2.1)
where data stands for the observed data, true stands for the true values of the kinematic variables describing the decay, $d\Phi_{n}$ are the phase space volume elements of both $\tau$ products, $\mathcal{M}(\text{true},m_{\text{test}})$ is matrix element responsible for the investigated process ($H\rightarrow\tau^{+}\tau^{-}$), $W(\text{data}|\text{true})$ is a transfer function between true and observed values of variables and $\mathcal{N}$ is arbitrary normalization factor. The $m_{\text{test}}$ is the hypothesis for the di-tau invariant mass, the likelihood we want to calculate.
@@ 2.1 The phase space
    $\tau^{-}\rightarrow\nu_{\tau}+\text{hadrons}$, denoted usually as all $\tau_{h}$.
Regardless of the decay channel, among the products we will always find neutrinos that are undetectable by the detector. To make the analysis clearer, we will introduce the notion of the visible and invisible four-momentum. The $(E^{\text{vis}},\overline{p}^{\text{vis}})$ will refer to the energy and momentum of electron, muon, or hadrons in the final states, while $(E^{\text{inv}},\overline{p}^{\text{inv}})$ will describe the sum of neutrinos’ four-momenta. We will also denote $x$ as the fraction of energy that is carried by the visible products of the decay:
@@
To fully describe the kinematics of a decaying tauon, we need three additional parameters that describe the kinematics of invisible products. We usually use $\theta_{\text{GJ}}$ – Gottfried-Jackson angle, which is an angle between the tau lepton momentum and sum of visible momenta; $\phi$ – azimuth angle describing $\tau$ momentum position on a cone centered around the visible momentum (Fig. 1) and $m_{\nu\nu}$, an invariant mass of the neutrino system. It is worth noting that in the case of hadron decay, $m_{\nu\nu}$ is equal to zero (assuming $m_{\nu}=0$), as there is only one neutrino in the final state.
Since taus originate from a heavy particle ($m\gg m_{\tau}$) decay, it is justified to assume that the tau decay products in the laboratory frame are collinear with the direction of the tau lepton momentum. Moreover, this assumption is supported by the poor resolution of the $\overline{E}_{T}^{\text{miss}}$ measurement. For light decay products, the maximal value of $\theta_{GJ}$ can be approximated as:
@@
where $\gamma$ is the Lorentz factor of the $\tau$ lepton in the laboratory frame.
For example, assuming that the energy from the Higgs decay is split equally between two $\tau$ leptons, we have $\gamma\approx 35$, and the $\theta_{GJ}$ is around $1.63^{\circ}$. Therefore, we assume:
@@
and leave the $\phi$ angle degenerate as a result.
Figure 1: Example of tau decay in electron channel $\tau^{-}\rightarrow e^{-}\nu_{\tau}\bar{\nu}_{e}$. The top picture shows the division of decay products – invisible products (neutrinos) are marked with red, while visible products are marked with blue. Light-grey color marks the $\tau^{-}$ decay cone. Bottom picture shows the definition of Gottfried-Jackson $\theta_{GJ}$ and the $\phi$ angles.
This translates to a physical interpretation in which both the visible and invisible systems are collinear with the direction of the $\tau$ lepton momentum, and:
@@
The range of the integration variables is constrained by the kinematics of tau lepton decay. For the $x$ it is:
@@
Detailed derivation of those limits can be found in the A. Finally, with the collinear approximation, we obtain a simple formula for the $\tau\tau$ invariant mass as a function of the invariant mass of the visible products, and the $x$ parameters:
@@
where $x_{1}$ and $x_{2}$ are the fractions of energy carried by the visible products of each $\tau$, respectively.
@@ 2.2 Matrix elements
The matrix element is maximally simplified to a uniform distribution over the phase space of the decay products:
$\left|\mathcal{M}(p,m_{\text{test}})\right|=\prod^{2}_{i=1}|\text{BW}^{(i)}_{\tau}|^{2},$ (2.10)
where we substitute the Breit-Wigner distribution with the narrow-width approximation:
$|\text{BW}^{(i)}_{\tau}|^{2}=\frac{\pi}{m_{\tau}\Gamma_{\tau}}\delta(m_{\tau}-x_{i}m_{i,\text{vis}}).$ (2.11)
In the likelihood calculation, we omit the normalization factor, as it is the same for all events and does not affect the position of the likelihood maximum. The likelihood is then given by phase space volume consistent with observed data, and values of the $x_{1,2}$ under the test:
$\begin{split}\mathcal{L}(m_{\text{test}}|\text{data})\sim\int d\Phi_{n}\,\left|\mathcal{M}(\text{true}|m_{\text{test}})\right|^{2}W(\text{data}|\text{true})=&\\ \int d\Phi_{n}\,\delta\left(m_{\text{test}}-\frac{m_{\text{vis}}}{\sqrt{x_{1}x_{2}}}\right)\delta(m_{\tau}-x_{1},\text{vis})\delta(m_{\tau}-x_{2},\text{vis})W(\text{data}|\text{true})\end{split}$ (2.12)
@@ 2.3 Transfer functions
We assume the $\overline{E}_{T}^{\text{miss}}$ uncertainty is much bigger than the uncertainty of the kinematic variables of the visible products of the $\tau$ decay. Therefore, we approximate the transfer functions for the visible products with identity. The $\overline{E}_{T}^{\text{miss}}$ transfer function is modelled with the normal distribution, with event-by-event covariance matrix as it was in the original cSVfit algorithm [4]. The covariance matrix of the reconstructed $\overline{E}_{T}^{\text{miss}}$ is used as an input of the algorithm. If such an event-by-event covariance matrix is not available, a constant covariance matrix can be used.
$W(\overline{E}_{T}^{\text{rec}}|\overline{E}_{T}^{\text{true}})=\mathcal{N}\exp\left(-\frac{1}{2}(\overline{E}_{T}^{\text{true}}-\overline{E}^{\text{rec}}_{T})^{T}{V^{-1}_{\text{MET}}}(\overline{E}_{T}^{\text{true}}-\overline{E}^{\text{rec}}_{T})\right),$ (2.13)
where $(\overline{E}_{T}^{\text{true}}-\overline{E}^{\text{rec}}_{T})$ is the difference between true and reconstructed $\overline{E}_{T}^{\text{miss}}$ and $V_{\text{MET}}$ is the covariance matrix. The $\overline{E}_{T}^{\text{true}}$ is calculated as the sum of the momenta of the invisible $\tau$ decays products, calculated from the $x_{1}$ and $x_{2}$ hypotheses and collinear approximation:
$\overline{E}_{T}^{\text{true}}=\frac{\overline{p}^{\text{vis}}_{1}}{x_{1}}+\frac{\overline{p}^{\text{vis}}_{2}}{x_{2}}.$ (2.14)
@@ 2.4 Final likelihood
Since the transfer function depends only on the $x_{1,2}$ values, the likelihood formula factorizes, and the integral over the phase space can be calculated independently of the transfer function or the event data:
$\begin{split}\mathcal{L}(m_{\text{test}}|\text{data})\sim W(\text{data}|\text{true})\int d\Phi_{n}\,\left|\mathcal{M}(\text{true}|m_{\text{test}})\right|^{2}\end{split}$ (2.15)
Hence, the integration over the phase space can be performed analytically. This step is crucial for improving the algorithm’s efficiency. The form of the integrand depends on the $\tau\tau$ decay channel. In the hadronic decay channel, where the integration over the neutrino invariant mass is not performed, we obtain:
@@
where $x_{\text{min}}=\max\left(x_{2,min},\left(\frac{m_{\text{vis}}}{m_{\text{test}}}\right)^{2}\right)$ and $x_{\text{max}}=\min\left(1,\left(\frac{m_{\text{vis}}}{m_{\text{test}}}\right)^{2}\frac{1}{x_{1,min}}\right)$. For the semi-leptonic decay, we need to consider integration over $m_{\nu\nu}$ for the leptonic $\tau$ decay:
@@
Finally, for the di-leptonic decay, we need to integrate over the $m_{\nu\nu}$ for both tauons:
@@
$\mathcal{L}(m_{\text{test}}|\text{data})=W(\overline{E}_{T}^{\text{rec}}|\overline{E}_{T}^{\text{true}})I$ (2.19)

### v3

@@ Abstract
The quality of the invariant mass reconstruction of the di-$\tau$ system is crucial for searches and analyses of di-$\tau$ resonances. Due to the presence of neutrinos in the final state, the $\tau\tau$ invariant mass cannot be calculated directly at hadron colliders, where the longitudinal momentum sum constraint cannot be applied. A number of approaches have been adopted to mitigate this issue. The most general approach uses Matrix Element (ME) integration for likelihood estimation, followed by invariant mass reconstruction as the value that maximizes the likelihood. However, this method has a significant computational cost due to the need for integration over the phase space of the decay products. We propose an algorithm that reduces the computational cost by two orders of magnitude while maintaining the invariant mass reconstruction resolution at a level comparable to that of the ME-based method. Moreover, we introduce additional features that allow the estimation of the uncertainty of the reconstructed mass and the kinematics of the initial $\tau$ leptons (e.g. their momenta).
@@ 1 Introduction
The $\tau\tau$ invariant mass cannot be calculated from the quantities measured by experiments at hadron colliders, as there are at least two neutrinos in the final state. The only available information is the total transverse momentum of all neutrinos, estimated from the missing transverse energy, denoted as $\overline{E}_{T}^{\text{miss}}$. Even though the momenta of the remaining neutral and charged decay products of the $\tau$ lepton (visible products, denoted $\tau_{\text{vis}}$) can be measured precisely, there are two main difficulties:
@@
    the incomplete description of the final-state kinematics;
@@
    the $\overline{E}_{T}^{\text{miss}}$ resolution, which is one of the least precisely reconstructed observables in a hadronic environment.
Two main approaches to mitigate these issues, adopted by the ATLAS [9] and CMS [10] Collaborations, are based on maximum-likelihood parameter estimation, $\mathcal{L}(m_{\text{test}}\,|\,\text{data})$, where $m_{\text{test}}$ is the tested $\tau\tau$ invariant mass hypothesis and “data” denotes all information available from the detector. However, the two collaborations define the likelihood in different ways.
In the case of the ATLAS Collaboration, the likelihood is based on the probability density functions (PDFs) of the distance between neutrinos and visible $\tau$ decay products [9]. Hypotheses for the neutrino directions are generated, and the one yielding the maximum likelihood value is chosen. The reconstructed neutrino directions, together with the measured visible momenta, allow for the calculation of the $\tau\tau$ invariant mass. The PDFs are obtained from a large sample of Monte Carlo simulations of $Z\rightarrow\tau\tau$ decays.
The CMS Collaboration adopted a matrix element–based technique in two versions: one using the full matrix element for $pp\rightarrow H\rightarrow\tau\tau\rightarrow X$ with all relevant normalization factors included (denoted Secondary Vertex Fit, SVfit), and a simplified version (classic Secondary Vertex Fit, cSVfit) [10]. In the simplified version, only terms corresponding to the $H\rightarrow\tau\tau\rightarrow X$ decay are retained. Both methods show similar performance in terms of mass resolution, but the simplified method requires significantly less computational time—around $0.25$ s on a modern computer. However, this is still too slow for complex analyses, where systematic uncertainty estimates require multiple invariant-mass calculations per event and the event samples to be processed are of the order of $10^{6}$ events.
In this article, we introduce a fast algorithm (fastMTT) inspired by cSVfit. We combine the idea of cSVfit with the well-known collinear approximation (CA) [11]. By applying the CA to the likelihood calculation, we achieve an approximately two orders of magnitude speedup in execution time while maintaining a comparable mass resolution.
The algorithm has been briefly described in the proceedings of the Epiphany Conference 2025 [12], where only the final analytical form of the method was presented. The present paper provides a complete and self-contained description of the algorithm, including a detailed motivation and derivation of the underlying formulae, the introduction of a regularization term, and additional features such as an uncertainty estimate and the reconstruction of the $\tau$-lepton kinematics. Moreover, the method’s performance is studied in greater detail and compared to existing approaches.
@@ 2 Maximum likelihood estimation
$\mathcal{L}(\text{data}|m_{\text{test}})=$ $\,\mathcal{N}\int d\Phi_{n}\,\left|\mathcal{M}(m_{\text{test}})\right|^{2}W(\text{data}|m_{\text{test}}),$ (2.1)
where data stands for the observed data, $d\Phi_{n}$ are the phase space volume elements of both $\tau$ products, $\mathcal{M}(m_{\text{test}})$ is the matrix element responsible for the investigated process ($H\rightarrow\tau^{+}\tau^{-}$), $W(\text{data}|m_{\text{test}})$ is a transfer function between true and observed values of variables and $\mathcal{N}$ is a normalization factor. In Eq. 2.1, $m_{\text{test}}$ denotes the tested di-$\tau$ invariant mass hypothesis, for which the likelihood is evaluated.
@@ 2.1 The phase space
    $\tau^{-}\rightarrow\nu_{\tau}+\text{hadrons}$, usually denoted as $\tau_{h}$.
Regardless of the decay channel, among the products we will always find neutrinos that are undetectable by the detector. To make the analysis clearer, we will introduce the notion of visible and invisible four-momentum. The $(E^{\text{vis}},\overline{p}^{\text{vis}})$ will refer to the four-vector of electrons, muons, or hadrons in the final states, while $(E^{\text{inv}},\overline{p}^{\text{inv}})$ will describe the sum of neutrinos’ four-momenta. We will also denote $x$ as the fraction of energy that is carried by the visible products of the decay:
@@
To fully describe the kinematics of a decaying tau, we need three additional parameters that describe the kinematics of invisible products. It is customary to use $\theta_{\text{GJ}}$ – Gottfried-Jackson angle [13] [14], which is the angle between the tau lepton momentum and sum of visible momenta in the tau rest frame; $\phi$ – azimuth angle describing the position of the $\tau$ momentum on a cone centered around the visible momentum (Fig. 1) and $m_{\nu\nu}$, an invariant mass of the neutrino system. It is worth noting that in the case of hadronic decay, $m_{\nu\nu}$ is equal to zero (assuming $m_{\nu}=0$), as there is only one neutrino in the final state.
Figure 1: Example of tau decay in the electron channel $\tau^{-}\rightarrow e^{-}\nu_{\tau}\bar{\nu}_{e}$. The top picture shows the division of decay products – invisible products (neutrinos) are marked with red, while visible products are marked with blue. Light-grey color marks the cone defined by the Gottfried-Jackson angle around $e^{-}$. Bottom picture shows the definition of Gottfried-Jackson $\theta_{GJ}$ and the $\phi$ angles.
Since taus originate from a heavy particle ($m\gg m_{\tau}$) decay, it is therefore reasonable to assume that the tau decay products in the laboratory frame are collinear (due to the Lorentz boost) with the direction of the tau lepton momentum. For light decay products, the maximal value of $\theta_{GJ}$ can be approximated as:
@@
where $\gamma$ is the Lorentz factor of the $\tau$ lepton in the laboratory frame. For example, assuming that the energy from the Higgs boson decay is split equally between two $\tau$ leptons, we have $\gamma\approx 35$, and $\theta_{GJ}$ is around $1.63^{\circ}$. Therefore, we assume
@@
and leave the $\phi$ angle degenerate as a result. This translates to a physical interpretation in which both the visible and invisible systems are collinear with the direction of the $\tau$ lepton momentum, and:
@@
The range of the integration variables is constrained by the kinematics of tau lepton decay. For $x$ it is:
@@
Detailed derivation of those limits can be found in the A.
Finally, with the collinear approximation, we obtain a simple formula for the $\tau\tau$ invariant mass as a function of the invariant mass of the visible products, and the $x$ parameters:
@@
where $x_{1}$ and $x_{2}$ are the fractions of energy carried by the visible products of each $\tau$, respectively, and $m_{\text{vis}}$ is the visible mass of the full di-$\tau$ system.
In the following, the likelihood is evaluated as a function of the parameters $(x1,x2)$, with the test mass $m_{\text{test}}$ understood as the derived quantity defined in Eq. (2.9).
@@ 2.2 Matrix elements
The matrix element is maximally simplified to include only the Breit–Wigner distribution of the $\tau$ lepton:
$\left|\mathcal{M}(p,m_{\text{test}})\right|^{2}=\prod_{i=1}^{2}\left|\mathrm{BW}^{(i)}_{\tau}\right|^{2},$ (2.10)
where $p$ denotes tau four-momentum.
The Breit–Wigner distribution is substituted using the narrow-width approximation:
$\left|\mathrm{BW}^{(i)}_{\tau}\right|^{2}=\frac{\pi}{m_{\tau}\Gamma_{\tau}}\,\delta\!\left(m_{\tau}^{2}-p^{2}_{\text{test}}\right).$ (2.11)
After integrating over the Dirac delta distribution, we are left solely with the phase space described in Section 2.1. The likelihood is then given by the phase-space volume consistent with the observed data and with the values of $x_{1,2}$ under the test mass:
$\begin{split}\mathcal{L}(\text{data}|x_{1},x_{2})\sim\int d\Phi_{n}\,\left|\mathcal{M}(m_{\text{test}}|x_{1},x_{2})\right|^{2}W(\text{data}|x_{1},x_{2})=&\\ \int d\Phi_{n}\,\delta\left(m_{\text{test}}-\frac{m_{\text{vis}}}{\sqrt{x_{1}x_{2}}}\right)W(\text{data}|x_{1},x_{2})\end{split}$ (2.12)
In the likelihood calculation, we omit the normalization factor, as it is the same for all events and therefore does not affect the position of the likelihood maximum.
@@ 2.3 Transfer functions
We assume the $\overline{E}_{T}^{\text{miss}}$ uncertainty is much bigger than the uncertainty of the kinematic variables of the visible products of the $\tau$ decay. Therefore, we approximate the transfer functions for the visible products with identity. The $\overline{E}_{T}^{\text{miss}}$ transfer function is modeled with the normal distribution, with event-by-event covariance matrix as it was in the original cSVfit algorithm [10]. The covariance matrix of the reconstructed $\overline{E}_{T}^{\text{miss}}$ is used as an input of the algorithm. If such an event-by-event covariance matrix is not available, a constant covariance matrix can be used.
$W(\overline{E}_{T}^{\text{rec}}|\overline{E}_{T}^{\text{test}})=\mathcal{N}\exp\left(-\frac{1}{2}(\overline{E}_{T}^{\text{test}}-\overline{E}^{\text{rec}}_{T})^{T}{V^{-1}_{\text{MET}}}(\overline{E}_{T}^{\text{test}}-\overline{E}^{\text{rec}}_{T})\right),$ (2.13)
where $(\overline{E}_{T}^{\text{test}}-\overline{E}^{\text{rec}}_{T})$ is the difference between the test and reconstructed missing transverse energy, and $V_{\text{MET}}$ is the reconstruction covariance matrix. The $\overline{E}_{T}^{\text{test}}$ is calculated as the sum of the momenta of the invisible $\tau$ decay products, calculated from the $x_{1}$ and $x_{2}$ hypotheses and collinear approximation:
$\overline{E}_{T}^{\text{test}}=\frac{\overline{p}^{\text{vis}}_{1}(1-x_{1})}{x_{1}}+\frac{\overline{p}^{\text{vis}}_{2}(1-x_{2})}{x_{2}}.$ (2.14)
@@ 2.4 Final likelihood
For the final step, we factorize the likelihood by taking $\overline{E}_{T}^{\text{test}}$ transfer function outside the integral:
$\begin{split}\mathcal{L}(\text{data}|x_{1},x_{2})\sim W(\text{data}|x_{1},x_{2})\int d\Phi_{n}\,\left|\mathcal{M}(m_{\text{test}}|x_{1},x_{2};m_{\text{vis}})\right|^{2}.\end{split}$ (2.15)
This factorization is motivated by two observations. First, each component independently peaks near the true invariant mass of the system; the corresponding validation studies are shown in B. Second, once separated, the transfer function and the phase–space term capture largely distinct pieces of information: the transfer function constrains the true missing transverse energy primarily, while the phase–space factor uses all remaining kinematic information of the event. Consequently, the residual correlation between them is weak.
Given two likelihood contributions that independently favor the true mass, using their product provides a practical and robust estimator of the overall likelihood profile.
With this factorization, the phase–space integral becomes analytically tractable, which is essential for achieving the computational performance required by the algorithm.
The form of the integrand depends on the $\tau\tau$ decay channel. In the hadronic decay channel, where the integration over the neutrino invariant mass is not performed, we obtain:
@@
where $x_{\text{min}}=\max\left(x_{2,min},\left(\frac{m_{\text{vis}}}{m_{\text{test}}}\right)^{2}\right)$ and $x_{\text{max}}=\min\left(1,\left(\frac{m_{\text{vis}}}{m_{\text{test}}}\right)^{2}\frac{1}{x_{1,min}}\right)$. For the semi-leptonic decay, we need also to consider integration over $m_{\nu\nu}$ for the leptonic $\tau$ decay:
@@
Finally, for the di-leptonic decay, we need to integrate over the $m_{\nu\nu}$ for both tau leptons:
@@
$\mathcal{L}(\text{data}|x_{1},x_{2})=W(\overline{E}_{T}^{\text{rec}}|\overline{E}_{T}^{\text{test}})I$ (2.19)